my brother actually needed help on this one, I forgot all of my Algebra I work. the problem was x^3 and the choices were a) x=2 b) x=5 and c) x=-3
I dont get it at all to explain it to him! help! thanks sooo much!
-Emily

Answer by mukhopadhyay(490)

I will help you memorizing algebra for your brother. You are a good sister though.
x^3 = (x)(x)(x)
1. When x=2, x^3 = 2*2*2 = 8
2. When x=5, x^3 = 5*5*5 = 125
3. When x=3, x^3 = 3*3*3 = 27